OpenStudy (anonymous):

solve -17≤ 4x-6≤ -5 the solution is

OpenStudy (anonymous):

x = -2, -1, 0

OpenStudy (anonymous):

-11/4<=x<=1/4

OpenStudy (anonymous):

-17≤ 4x-6≤ -5 ( -17+6)/4≤ x ≤ (-5+6)/4 \[-\frac{11}{4}\le x \le \frac{1}{4}\]